The transition temperature of a BCS superconductor with an average atomic mass of 199.5 is 4.18K. What is the transition
temperature for an average atomic mass of 203.4?

A) 4.185
B) 4.145
C) 4.226
D) 4.105
1
Expert's answer
2021-10-14T11:34:38-0400
"T_1\sim\frac 1{\sqrt{m_{a1}}},\\\space\\
T_2\sim\frac 1{\sqrt{m_{a2}}},\\\space\\
\frac{T_1}{T_2}=\sqrt{\frac{m_{a2}}{m_{a1}}},\\\space\\
T_2=T_1\sqrt{\frac{m_{a1}}{m_{a2}}}=4.14\text{ K}."